Solve the following equations by completing squares and then check their answers.
a^2 - a = 6

Found 2 solutions by josgarithmetic, ewatrrr:
Answer by josgarithmetic(39617)
Answer by ewatrrr(24785)

a^2 - a = 6
(a-.5)^2 = 6 + .25
(a-.5)^2 = 6.25
a^2 - a + .25 = 6 + .25